A bicyclist rode up a mountain road at 12 km/h and then back down 30 km/h. If the round trip took 3.5 h, how long did the ride up the mountain take?
Answer by funmath(2873)

A bicyclist rode up a mountain road at 12 km/h and then back down 30 km/h. If the round trip took 3.5 h, how long did the ride up the mountain take?
distance (d)=rate (r)*time(t) -->> d=rt
Let the time (t) up the mountain be: t
Then the rest of the 3.5 hrs is the time to get down: (3.5-t)
Then the distance up is: 12t
and the distance down is: 30(3.5-t)
The distance up a mountain is the same as the distance down so:
12t=30(3.5-t)
12t=105-30t
12t+30t=105-30t+30t
42t=105
42t/42=105/42
t=2.5
***The trip up the mountain took: t=2.5 hrs
You weren't asked, but you can also find out the time for the trip down the mountain, just for fun. You can even find out how long the trip up was, or the trip down, or the round trip. Play around with this until you feel comfortable.
Happy Calculating!!!